# Smallest Suffixient Sets as a Repetitiveness Measure

### v1

@@ 4 Sensitivity of $\chi$ to String Operations
For Claim 2, note that exactly $k$ substrings of length $k$ are lost when substituting the last b of $w$ by c: those of the form $\texttt{b}^{i}\texttt{a}^{k-i}$ with $i>0$. This means that substrings ending in $\texttt{b}^{i}\texttt{a}^{k-i-1}$ with $0<i<k$ are not right-maximal in $w^{\prime}$, hence, $2(k-1)$ supermaximal extensions are lost. Moreover, $\texttt{b}^{k-2}\texttt{c}$ is a new supermaximal extension, but its pair $\texttt{b}^{k-1}$ is a suffix of $\texttt{a}\texttt{b}^{k-1}$. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)+1$.
@@
For Claim 3, the analysis is similar to Claim 2, but in $w^{\prime}$, $\texttt{b}^{k-1}$ remains as a supermaximal extension, so $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)+2$. For Claim 4, the analysis is similar to Claim 1, but in $w^{\prime}$, $\texttt{a}^{k-2}\texttt{b}$ b appears, while $\texttt{a}^{k-1}\texttt{b}$ does not. ∎
@@
Let $k>0$. Let $w_{k}=\prod_{i=0}^{k-1}\texttt{c}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{i}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\$}_{i}$ on the alphabet $\Sigma=\{\texttt{a},\texttt{b},\texttt{c}\}\cup\bigcup_{i\in[0\mathinner{.\,.}k-1]}\{\text{\tt\#}_{i},\text{\tt\$}_{i}\}$. It holds $\mathtt{sre}(w_{k})=5k-1$ and $\mathtt{sre}(w_{k}^{R})=4k$.
@@
Note that, by construction, any substring of $w_{k}$ containing #$\text{\tt\#}_{i}$ or \$$\text{\tt\$}_{i}$ is not right-maximal. We list the supermaximal extensions of $w_{k}$:
@@
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{i}$ and $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\$}_{i}$ for $i\in[0\mathinner{.\,.}k-1]$,
@@
This sums to a total of $15k-$ 1 supermaximal extensions in $w_{k}$. In the reversed string $w_{k}^{R}=\prod_{i=0}^{k-1}\text{\tt\$}_{k-i-1}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{k-i-1}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\texttt{c}$, we have instead:
@@
    $\texttt{b}\texttt{a}^{k-1}$, \$$\text{\tt\$}_{k-1}$ and $\texttt{a}^{k-2}\texttt{c}\text{\tt\$}_{k-2}$
@@
    $\texttt{a}^{k-i-1}\texttt{c}\text{\tt\$}_{k-i-2}$ for $i\in[1\mathinner{.\,.}k-2]$,
@@
This sums to a total of $4k$ supermaximal extensions in $w_{k}^{R}$, of length $|w_{k}|=|w_{k}^{R}|=k(2k+3)$. Thus, $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=k-1=\Theta(\sqrt{n})$. ∎
Finally, we show upper bounds on the sensitivity of $\mathtt{sre}$ to string operations. [^2] For the multiplicative sensitivity, we assume $w$ and $w^{\prime}$ are not unary strings, as otherwise $\mathtt{sre}(w)$ or $\mathtt{sre}(w^{\prime})$ would be 00. This does not happen with $\chi$.
@@
$\mathtt{sre}(w^{\prime})-\mathtt{sre}(w)=O\left(\delta\max\left(1,\log(n/\delta\log\delta)\right)\log\delta\right)\text{ and }$
$\mathtt{sre}(w^{\prime})\hskip 3.0pt/\hskip 4.0pt\mathtt{sre}(w)=O\left(\max\left(1,\log(n/\delta\log\delta)\right)\log\delta\right).$
@@
It follows because the multiplicative sensitivity of $\delta$ to the string operations considered and reversals is $O(1)$ [1], the known relations $\delta\leq\gamma\leq\chi\leq 2\overline{r}$ [4], and the upper bound $r=O(\delta\max(1,\log(n/\delta\log\delta))\log\delta)$ [17].∎
@@ 5.1 Proving $\chi=O(r)$
We first prove that $\chi$ is asymptotically upper-bounded by the number $r$ of runs in the BWT of the sequence.
@@
Let $xa$ be a super-maximal right-extension in $w\text{\tt\$}$. We distinguish, in the BWT-matrix of $w\text{\tt\$}$, the sets $\mathcal{R}_{xc}$ of rotations starting with $xc$ where $c\in\Sigma\cup\{\text{\tt\$}\}$. Because $x$ is right-maximal, at least 2 of these blocks are non-empty; i) the set $\mathcal{R}_{xa}$; ii) some set $\mathcal{R}_{xb}$ of rotations starting with $xb$, where $b\neq a$.
We claim that the last characters of the rotations in $\mathcal{R}_{xa}$ must be disjoint from the last characters of rotations in $\mathcal{R}_{xb}$, for any $b\neq a$. Suppose by contradiction that there are two rotations of $w\text{\tt\$}$, of the form $xa\dots c$ and $xb\dots c$. Then, $cxa$ and $cxb$ are circular substrings of $w\text{\tt\$}$. Note that $cx$ does not contain \$, otherwise, as \$ is unique and both circular substrings have the same length, $cxa$ and $cxb$ would have to be the same string, yet $a$ and $b$ are different. This implies $cxa$ and $cxb$ have to be substrings of $w\text{\tt\$}$. Moreover, $cx$ is a right-maximal substring of $T$, and $cxa$ is a one-character right-extension that contains $xa$ as a suffix, contradicting that $xa$ is super-maximal.
Let $a_{1}<\cdots<a_{t}$ be all the characters such that $xa_{i}$ is a super-maximal right-extension, for $t\geq 2$. Those induce (not necessarily consecutive) BWT areas $\mathcal{R}_{xa_{i}}$. The argument of the previous paragraph applies to any pair $a=a_{i}$ and $b=a_{i+1}$, and implies that a new BWT run must start between the first row following $\mathcal{R}_{xa_{i}}$ and the first row of $\mathcal{R}_{xa_{i+1}}$, for all $i=2,\ldots,t$. The string $x$ then induces $t$ super-maximal right-extensions and $t-1$ BWT runs. The worst ratio between contributions to $\mathtt{sre}$ and to $r$ is 2 to 1, which occurs when $t=2$.
The contributions can be summed for other super-maximal right-extensions $ya^{\prime}_{i}$ if $y$ does not prefix or is prefixed by $x$, because the corresponding BWT ranges are disjoint. We now focus in the case where $y$ is a prefix of $x$. Since $y$ is right-maximal, its range $\mathcal{R}_{y}$ strictly contains $\mathcal{R}_{x}$. Further, $\mathcal{R}_{x}$ is completely contained in the range of one of the extensions of $y$, precisely $\mathcal{R}_{ye}$ where $e=x[|y|+1]$. The characters $a^{\prime}_{1}<\cdots<a^{\prime}_{t^{\prime}}$ such that $ya^{\prime}_{i}$ is a super-maximal right-extension also induce $t^{\prime}-1$ BWT runs, starting between the row following the areas $\mathcal{R}_{ya^{\prime}_{i}}$ and the first row of the areas $\mathcal{R}_{ya^{\prime}_{i+1}}$. Importantly, those induced run start positions are disjoint from those induced by $x$: the first run start position induced by $x$ is after $\mathcal{R}_{xa_{1}}$, whereas the only run start induced by $y$ inside $\mathcal{R}_{ye}$ is at the first row of $\mathcal{R}_{xa_{1}}$ or earlier. Therefore, the induced runs of $x$ and $y$ are disjoint too.
Adding over all super-maximal extensions, it follows that $r\geq\chi/2$. ∎
@@ 5.2 A family with $\chi=o(v)$ (and thus $o(r)$)
We will now show that $\chi=o(v)$ on the so-called Fibonacci words, which also implies $\chi=o(r)$ in that string family because $v=O(r)$ [27]. Combined with Lemma 9, this implies that $\chi$ is a strictly smaller measure than $r$. Instead, $\chi$ is incomparable with $v$, as we show later. In our way, we obtain some relevant byproducts about the structure of suffixient sets on Fibonacci, and more generally, episturmian words.
@@
Let w be episturmian but not epistandard. There exists some epistandard word s with the same set of substrings, i.e., $\mathcal{F}_{\textbf{w}}=\mathcal{F}_{\textbf{s}}$ [8]. Therefore, for any episturmian word w, there exist exactly $\sigma$ suffix-chains of right-extensions.
It follows that for any substring $\textbf{w}[i\mathinner{.\,.}j]$ of any episturmian word w, $\mathtt{sre}\leq\sigma$, and the supermaximal extension for each $a\in\Sigma$ appearing in $\textbf{w}[i\mathinner{.\,.}j]$ is the longest reversed prefix of w followed by $a$ appearing in $\textbf{w}[i\mathinner{.\,.}j]$, and having another occurrence within $\textbf{w}[i\mathinner{.\,.}j]$ followed by another character. ∎
Combining this result with Lemma 2, we have the following bound.
@@
Every Fibonacci word $F_{k}\text{\tt\$}$ has a suffixient set of size at most 4. For $k\geq 6$, the only smallest suffixient sets for $F_{k}\text{\tt\$}$ are $\{f_{k}+1,f_{k}-1,f_{k-1}-1,p\}$, where $p\in\{f_{k-2}+1,2f_{k-2}+1\}$.
@@
Note that $F_{k-1}=H_{k-1}dc$ prefixes $F_{k}$. On the other hand, we can write $F_{k}=F_{k-1}F_{k-2}=F_{k-2}F_{k-3}F_{k-2}=F_{k-2}F_{k-1}^{\prime}=F_{k-2}H_{k-1}cd$. Therefore, string $H_{k-1}$ is right-maximal in $F_{k}$. Its extensions, $H_{k-1}d$ and $H_{k-1}c$, are super-maximal because there are no other occurrences of $H_{k-1}$ in $F_{k}$: (i) $H_{k-1}$ cannot occur starting at positions $f_{k-2}+2$ or $f_{k-2}+3$ because it occurs at $f_{k-2}+1$, so $H_{k-1}$ should match itself with an offset of 1 or 2, which is impossible because it prefixes $F_{k-1}$ and all $F_{k-1}$ for $k-1\geq 5$ start with abaab; (ii) $H_{k-1}$ cannot occur starting at positions $2$ to $f_{k-2}$ because its prefix $F_{k-2}$ should occur inside the prefix $F_{k-2}F_{k-2}$ of $F_{k}=F_{k-2}F_{k-1}^{\prime}=F_{k-2}F_{k-2}F_{k-3}^{\prime}$, and so $F_{k-2}$ should equal a rotation of it, which is impossible [7, Cor. 3.2]. The two positions following $H_{k-1}$, $f_{k-1}-1$ and $f_{k}-1$, then appear in any suffixient set.
On the other hand, $F_{k-2}$ is followed by \$ in $F_{k}\text{\tt\$}$, and it also prefixes $F_{k}=F_{k-2}F_{k-1}^{\prime}$, therefore $F_{k-2}$ is right-maximal. The first occurrence is preceded by $F_{k-1}$, and hence by $c$, and the second by no symbol. $F_{k-2}$ also occurs in $F_{k}$ at position $f_{k-2}+1$, as seen above, preceded by $F_{k-2}$ and thus by $d$. There are no other occurrences of $F_{k-2}$ in $F_{k}$ because (i) it cannot occur starting at positions $2$ to $f_{k-2}$ by the same reason as point (ii) of the previous paragraph; (ii) it cannot appear starting at positions $f_{k-2}+2$ to $f_{k-1}-2$ because $F_{k}=F_{k-2}F_{k-2}F_{k-3}^{\prime}$ and $F_{k-3}^{\prime}[1,f_{k-3}-2]=F_{k-3}[1,f_{k-3}-2]=F_{k-2}[1..f_{k-3}-2]$, thus such an occurrence would also match a rotation of $F_{k-2}$, which is impossible as noted above; (iii) it cannot appear starting at positions $f_{k-1}-1$ or $f_{k-1}$ because, since it matches at position $f_{k-1}+1$, $F_{k-2}$ would match itself with an offset of 1 or 2, which is impossible as noted in point (i) of the previous paragraph. The right-extensions of $F_{k-2}$ are then super-maximal. The one followed by \$ occurs ending at position $f_{k}+1$. The other two are followed by a because they are followed by $F_{k-2}$ and by $F_{k-3}^{\prime}$ and all $F_{k}$ for $k\geq 2$ start with a. We can then choose either ending position for a suffixient set, $f_{k-2}+1$ or $12f_{k-2}+$ 1. ∎
@@
It follows from Lemma 11 and the fact that $v=\Omega(\log n)$ on the odd Fibonacci words [27, Thm. 28].

### v2

@@ 4 Sensitivity of $\chi$ to String Operations
For Claim 2, note that exactly $k$ substrings of length $k$ are lost when substituting the last b of $w$ by c: those of the form $\texttt{b}^{i}\texttt{a}^{k-i}$ with $i>0$. This means that substrings ending in $\texttt{b}^{i}\texttt{a}^{k-i-1}$ with $0<i<k$ are not right-maximal in $w^{\prime}$, hence, 2(k1)2(k-1) super-maximal extensions are lost. Moreover, $\texttt{b}^{k-2}$ is still a right-maximal substring, since $\texttt{b}^{k-1}$ and $\texttt{b}^{k-2}\texttt{c}$ occur in $w^{\prime}$. Observe that only $\texttt{b}^{k-2}\texttt{c}$ is a super-maximal extension, while $\texttt{b}^{k-1}$ is a suffix of $\texttt{a}\texttt{b}^{k-1}$. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)+1$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-3$.
@@
For Claim 3, the analysis is similar to Claim 2, but in $w^{\prime}$, $\texttt{b}^{k-1}$ remains as a super-maximal extension. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)+2$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-4$.
For Claim 4, the analysis is similar to Claim 1, but in $w^{\prime}$, $\texttt{b}\texttt{a}^{k-2}\texttt{b}$ appears, while $\texttt{a}^{k-1}\texttt{b}$ does not. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-2$. ∎
@@
Let $k>0$. Let $w_{k}=\prod_{i=0}^{k-1}\texttt{c}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{i}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\textdollar}_{i}$ on the alphabet $\Sigma=\{\texttt{a},\texttt{b},\texttt{c}\}\cup\bigcup_{i\in[0\mathinner{.\,.}k-1]}\{\text{\tt\#}_{i},\text{\tt\textdollar}_{i}\}$. It holds $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=k-1$.
@@
Observe that by construction, any substring of $w_{k}$ containing $\text{\tt\#}_{i}$ or $\text{\tt\textdollar}_{i}$ is not right-maximal, as these symbols are unique. Hence, the right-extensions of $w_{k}$ cannot cross from one side to the other side of those special delimiters. Moreover, substrings of the form $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[0\mathinner{.\,.}k-1]$ appear exactly twice in $w_{k}$ and their right-extensions are super-maximal. By looking at the structure of the string $w_{k}$ and carefully analyzing its right-extensions, one can verify that the super-maximal right-extensions of $w_{k}$ are the following:
@@
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{i}$ and $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\textdollar}_{i}$ for $i\in[0\mathinner{.\,.}k-1]$,
@@
This sums to a total of 5k15k-1 super-maximal extensions in $w_{k}$. In the reversed string $w_{k}^{R}=\prod_{i=0}^{k-1}\text{\tt\textdollar}_{k-i-1}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{k-i-1}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\texttt{c}$, we have instead:
@@
    $\texttt{b}\texttt{a}^{k-1}$ and $\text{\tt\textdollar}_{k-1}$,
@@
    $\texttt{a}^{k-i-1}\texttt{c}\text{\tt\textdollar}_{k-i-2}$ for $i\in[1\mathinner{.\,.}k-2]$, and $\texttt{a}^{k-2}\texttt{c}\text{\tt\textdollar}_{k-2}$,
@@
This sums to a total of $4k$ super-maximal extensions in $w_{k}^{R}$. Thus, $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=(5k-1)-4k=k-1$. ∎
We give an example of the words $w_{k}$ and $w_{k}^{R}$ of Lemma 7, and their super-maximal right-extensions.
Example 1
Let $w_{3}=\texttt{c}\texttt{baa}\text{\tt\#}_{0}\texttt{baa}\text{\tt\textdollar}_{0}\texttt{c}\texttt{aba}\text{\tt\#}_{1}\texttt{aba}\text{\tt\textdollar}_{1}\texttt{c}\texttt{aab}\text{\tt\#}_{2}\texttt{aab}\text{\tt\textdollar}_{2}$. It can be verified that the super-maximal right-extensions of $w_{3}$ are:
  1. 1.
    baa and c;
  2. 2.
    $\texttt{baa}\text{\tt\#}_{0}$ and $\texttt{baa}\text{\tt\textdollar}_{0}$; $\texttt{aba}\text{\tt\#}_{1}$ and $\texttt{aba}\text{\tt\textdollar}_{1}$; $\texttt{aab}\text{\tt\#}_{2}$ and $\texttt{aab}\text{\tt\textdollar}_{2}$;
  3. 3.
    ca and cb; caa and cab;
  4. 4.
    aba and aab.
Similarly, let $w_{3}^{R}=\text{\tt\textdollar}_{2}\texttt{baa}\text{\tt\#}_{2}\texttt{baa}\texttt{c}\text{\tt\textdollar}_{1}\texttt{aba}\text{\tt\#}_{1}\texttt{aba}\texttt{c}\text{\tt\textdollar}_{0}\texttt{aab}\text{\tt\#}_{0}\texttt{aab}\texttt{c}$. The super-maximal right-extensions of $w_{3}^{R}$ are:
  1. 1.
    baa and $\text{\tt\textdollar}_{2}$;
  2. 2.
    $\texttt{baa}\text{\tt\#}_{2}$ and baac; $\texttt{aba}\text{\tt\#}_{1}$ and abac; $\texttt{aab}\text{\tt\#}_{0}$ and aabc;
  3. 3.
    $\texttt{ac}\text{\tt\textdollar}_{0}$; $\texttt{ac}\text{\tt\textdollar}_{1}$;
  4. 4.
    aba and aab.
One can see that $\mathtt{sre}(w_{3})=14$, $\mathtt{sre}(w_{3}^{R})=12$, and hence, $\mathtt{sre}(w_{3})-\mathtt{sre}(w_{3}^{R})=2$, as stated in Lemma 7.
Formally, the additive sensitivity of a measure of repetitiveness $\mu$ to a string operation $\rho$ can be defined as a function $AS_{\mu,\rho}:\mathbb{Z}^{+}\rightarrow\mathbb{R}$, where $AS_{\mu,\rho}(n)=\max_{w\in\Sigma^{n}}(\max_{w^{\prime}\in\rho(w)}(\mu(w^{\prime}))-\mu(w))$, that is the maximum achievable difference among all the strings. Overall, we obtain the following result on the additive sensitivity of $\mathtt{sre}$, which, by Corollary 1, can be written in terms of $\chi$.
Corollary 2
The following bounds on the additive sensitivity of the measure $\chi$ to string operations hold:
  1. 1.
    $AS_{\chi,\rho}(n)=\Omega(\log n)$ for $\rho\in\{\mathtt{ins},\mathtt{del},\mathtt{sub},\mathcal{R}(\cdot)\}$;
  2. 2.
    $AS_{\chi,\mathtt{rev}}(n)=\Omega(\sqrt{n})$, where $\mathtt{rev}(w)=\{w^{R}\}$.
Proof
Claim 1 follows by Lemma 6, where $n=|w|=\Theta(2^{k})$ and $AS_{\chi,\rho}(n)=\Omega(k)=\Omega(\log n)$, for all $\rho\in\{\mathtt{ins},\mathtt{del},\mathtt{sub},\mathcal{R}(\cdot)\}$. Claim 2 follows by Lemma 7, where $n=|w_{k}|=\Theta(k^{2})$ and $AS_{\chi,\mathtt{rev}}(n)=\Omega(k)=\Omega(\sqrt{n})$. ∎
Finally, we show upper bounds on the sensitivity of $\chi$ to string operations.
@@
$\chi(w^{\prime})-\chi(w)=O\left(\delta\max\left(1,\log(n/\delta\log\delta)\right)\log\delta\right)\text{ and }$
$\chi(w^{\prime})\hskip 3.0pt/\hskip 4.0pt\chi(w)=O\left(\max\left(1,\log(n/\delta\log\delta)\right)\log\delta\right).$
@@
To prove our thesis, we rely on the relations $\delta\leq\chi\leq 2\overline{r}$ [4] and $r=O(\delta\max(1,\log(n/\delta\log\delta))\log\delta)$ [17]. Moreover, since the multiplicative sensitivity of the measure $\delta$ to any of the string operations is $O(1)$ [1], for any $w\in\Sigma^{*}$ it holds $\overline{r}(w)=r(w^{R})=O(\delta\max(1,\log(n/\delta\log\delta))\log\delta)$. The thesis follows by considering the worst case, that is $\chi(w)=\Theta(\delta)$ and $\chi(w^{\prime})=\Theta(\delta\max(1,\log(n/\delta\log\delta))\log\delta)$.∎
@@ 5.1 Proving $\chi=O(r)$
We first prove that $\chi$ is asymptotically upper-bounded by the number $r$ of runs in the BWT of the sequence. As for the measure $\chi$, we assume that the BWT is computed after appending the \$ symbol.
@@
Let $x_{i}$ denotes the $i$th rotation of $w\text{\tt\textdollar}$ in lexicographic order, for each $i\in[1\mathinner{.\,.}|w|+1]$, and let $u_{i}$ be the longest common prefix between the rotations $x_{i},x_{i+1}$, for each $i\in[1\mathinner{.\,.}|w|]$. We further define $s:[1\mathinner{.\,.}n+1]\rightarrow[0\mathinner{.\,.}n]$ as $s(i)=j$ if $x_{i}=w[j+1\mathinner{.\,.}|w|]\text{\tt\textdollar}w[1\mathinner{.\,.}j]$, i.e., the number of cyclic shift to the right required to transform $x_{i}$ into $w\text{\tt\textdollar}$. [^2] The function $s$ mimics the well-known Suffix Array [23], here omitted for simplicity of exposition. As the symbol \$ occurs only once in $w\text{\tt\textdollar}$, the function $s$ is bijective.
Note that each right-extension of $w\text{\tt\textdollar}$ can be written as $u_{i}c$, for some $i\in[1\mathinner{.\,.}|w|]$ and $c\in\Sigma$. Consider now the set
$S=\bigcup_{i\in[1\mathinner{.\,.}|w|]}\{s(i)+|u_{i}|+1,s(i+1)+|u_{i}|+1\},$
that is the set of positions where the occurrences of the right-extensions $u_{i}c_{1}$ and $u_{i}c_{2}$ end in $w\text{\tt\textdollar}$, where $u_{i}c_{1}$ and $u_{i}c_{2}$ are the prefix of $x_{i}$ and $x_{i+1}$ respectively, for some $c_{1},c_{2}\in\Sigma$ such that $c_{1}<c_{2}$. It follows by construction that the set $S$ is a suffixient set of $w\text{\tt\textdollar}$.
We now show that $|S|\leq 2r$. Let us factorize each pair of consecutive rotations in the BWT-matrix as $x_{i}=u_{i}v_{i}c_{i}$ and $x_{i+1}=u_{i}v^{\prime}_{i}c_{i+1}$. Observe that $v_{i},v^{\prime}_{i}\neq\varepsilon$ [9, Corollary 8], $v_{i}[1]\neq v^{\prime}_{i}[1]$, and $c_{i}=\mathtt{BWT}(w\text{\tt\textdollar})[i]$ for all $i\in[1\mathinner{.\,.}|w|+1]$. A well-known property of the BWT-matrix is that if $c_{i}=c_{i+1}=c\in\Sigma$, then there exists $j\in[1\mathinner{.\,.}|w|]$ such that $x_{j}=cu_{i}v_{i}$ and $x_{j+1}=cu_{i}v^{\prime}_{i}$ [3]. As a consequence, one has that $s(j)+|u_{j}|+1=(s(i)-1)+(|u_{i}|+1)+1=s(i)+|u_{i}|+1$ and $s(j+1)+|u_{j}|+1=(s(i+1)-1)+(|u_{i}|+1)+1=s(i+1)+|u_{i}|+1$, and the procedure can be reiterated as long as $x_{j}$ and $x_{j+1}$ end with the same symbol. It follows that the same set can be written as
$S=\{s(i)+|u_{i}|+1,s(i+1)+|u_{i}|+1\mid i\in[1\mathinner{.\,.}|w|]\wedge\mathtt{BWT}[i]\neq\mathtt{BWT}[i+1]\},$
[…]
@@ 5.2 A Family with $\chi=o(v)$ (and thus $o(r)$)
We will now show that $\chi=o(v)$ on the so-called Fibonacci words, which also implies $\chi=o(r)$ in that string family because $v=O(r)$ [28]. Combined with Lemma 9, this implies that $\chi$ is a strictly smaller measure than $r$. In contrast, $\chi$ is incomparable with $v$, as we show later. On our way, we obtain some relevant byproducts about the structure of suffixient sets on Fibonacci, and more generally, episturmian words.
@@
Let w be episturmian but not necessarily epistandard. There exists some epistandard word s with the same set of substrings, i.e., $\mathcal{F}_{\textbf{w}}=\mathcal{F}_{\textbf{s}}$ [8]. Therefore, for any episturmian word w, there exist exactly $\sigma$ suffix-chains of right-extensions.
When considering substrings of w, the super-maximal right-extension in $\textbf{w}[i\mathinner{.\,.}j]$ ending with $a\in\Sigma$ is the longest right-extension of w ending with $a$ that remains a right-extension in $\textbf{w}[i\mathinner{.\,.}j]$. It follows that for any substring $\textbf{w}[i\mathinner{.\,.}j]$ of any episturmian word w, $\mathtt{sre}\leq\sigma$. ∎
Combining this result with Corollary 1, we obtain the following bound.
@@
Every Fibonacci word $F_{k}\text{\tt\textdollar}$ has a suffixient set of size at most 4. For $k\geq 6$, the only smallest suffixient sets for $F_{k}\text{\tt\textdollar}$ are $\{f_{k}+1,f_{k}-1,f_{k-1}-1,p\}$, where $p\in\{f_{k-2}+1,2f_{k-2}+1\}$.
@@
Note that $F_{k-1}=H_{k-1}dc$ prefixes $F_{k}$. On the other hand, we can write $F_{k}=F_{k-1}F_{k-2}=F_{k-2}F_{k-3}F_{k-2}=F_{k-2}F_{k-1}^{\prime}=F_{k-2}H_{k-1}cd$. Therefore, string $H_{k-1}$ is right-maximal in $F_{k}$. Its extensions, $H_{k-1}d$ and $H_{k-1}c$, are super-maximal because there are no other occurrences of $H_{k-1}$ in $F_{k}$: (i) $H_{k-1}$ cannot occur starting at positions $f_{k-2}+2$ or $f_{k-2}+3$ because it occurs at $f_{k-2}+1$, so $H_{k-1}$ should match itself with an offset of 1 or 2, which is impossible because it prefixes $F_{k-1}$ and all $F_{k-1}$ for $k-1\geq 5$ start with abaab; (ii) $H_{k-1}$ cannot occur starting at positions $2$ to $f_{k-2}$ because its prefix $F_{k-2}$ should occur inside the prefix $F_{k-2}F_{k-2}$ of $F_{k}=F_{k-2}F_{k-1}^{\prime}=F_{k-2}F_{k-2}F_{k-3}^{\prime}$, and so $F_{k-2}$ should equal a rotation of itself, which is impossible [7, Cor. 3.2]. The two positions following $H_{k-1}$, $f_{k-1}-1$ and $f_{k}-1$, then appear in any suffixient set.
On the other hand, $F_{k-2}$ is followed by \$ in $F_{k}\text{\tt\textdollar}$, and it also prefixes $F_{k}=F_{k-2}F_{k-1}^{\prime}$, therefore $F_{k-2}$ is right-maximal. The first occurrence is preceded by $F_{k-1}$, and hence by $c$, and the second by no symbol. $F_{k-2}$ also occurs in $F_{k}$ at position $f_{k-2}+1$, as seen above, preceded by $F_{k-2}$ and thus by $d$. There are no other occurrences of $F_{k-2}$ in $F_{k}$ because (i) it cannot occur starting at positions $2$ to $f_{k-2}$ by the same reason as point (ii) of the previous paragraph; (ii) it cannot appear starting at positions $f_{k-2}+2$ to $f_{k-1}-2$ because $F_{k}=F_{k-2}F_{k-2}F_{k-3}^{\prime}$ and $F_{k-3}^{\prime}[1,f_{k-3}-2]=F_{k-3}[1,f_{k-3}-2]=F_{k-2}[1..f_{k-3}-2]$, thus such an occurrence would also match a rotation of $F_{k-2}$, which is impossible as noted above; (iii) it cannot appear starting at positions $f_{k-1}-1$ or $f_{k-1}$ because, since it matches at position $f_{k-1}+1$, $F_{k-2}$ would match itself with an offset of 1 or 2, which is impossible as noted in point (i) of the previous paragraph. The right-extensions of $F_{k-2}$ are then super-maximal. The one followed by \$ occurs ending at position $f_{k}+1$. The other two are followed by a because they are followed by $F_{k-2}$ and by $F_{k-3}^{\prime}$ and all $F_{k}$ for $k\geq 2$ start with a. We can then choose either ending position for a suffixient set, $f_{k-2}+1$ or $2f_{k-2}+1$. ∎
@@
It follows from Lemma 11 and the fact that $v=\Omega(\log n)$ on the odd Fibonacci words [28, Thm. 28]. ∎